# MFC 5.0: An exascale many-physics flow solver

### v3

@@ 4.2.2 WENO-Z, TENO, and high-order non-uniform reconstruction
The WENO-Z scheme [68] enhances classical WENO methods by improving their spectral properties while being faster than WENO-M. The $(2k-1)$-th order WENO-Z weight $\omega_{r}^{(\mathrm{Z})}$ for the $r$-th stencil is computed as
@@
The Targeted Essentially Non-Oscillatory (TENO) scheme further reduces numerical dissipation by employing a stencil selection process similar to the ENO method [69]. A $K$-th order accurate TENO scheme uses $K-2$ candidate stencils of increasing width. Its equations are given by
@@
where $C_{T}$ is a smoothness threshold for stencil activation and $\tau$ is defined as in the WENO-Z scheme.
The explicit expressions for the seventh-order WENO reconstruction in our code are omitted here due to their considerable length. Instead, we present a generalized approach for reconstructing a $(2k-1)$-order scheme from cell averages on non-uniform grids. For each stencil, the reconstruction polynomial is obtained using the Lagrange interpolation method [70]:
@@
Here, $x_{j}$ denotes cell-center location, $x_{j+1/2}$ the cell-boundary location, and $\Delta x_{j}:=x_{j+1/2}-x_{j-1/2}$ is the cell width. For each stencil $r=0,1,\dots,k-1$, the $k$-cell stencil spans the interval from $x_{-1/2}$ to $x_{k-1/2}$. Consequently, the overall $(2k-1)$-cell stencil is centered at $x_{k-r-1}$, and the right flux is $p_{r}:=p(x_{k-r-1/2})$.
@@
where $\delta_{jm}$ is the Kronecker delta. The coefficients $\hat{b}_{r,j,m}$ for each cross term $\Delta f_{j}\Delta f_{m}$ and stencil $r$ are precomputed. Reformulating in terms of $\Delta f$ reduces the number of terms from $k$ to $k-1$ for $p_{r}$, and from $k(k+1)/2$ to $(k-1)k/2$ for $\beta_{r}$, thereby reducing the computational cost.
@@ 4.2.3 Strang splitting for stiff sub-grid dynamics
Sub-grid bubbles often exhibit rapid dynamics, necessitating a small time step for stable computation of the bubble source term. When the time scale of the background flow significantly exceeds that of the sub-grid bubbles, fully resolving the temporal dynamics becomes computationally prohibitive. To address this, we implement the Strang splitting scheme, which separates the time integration of the background flow and the sub-grid bubbles [71]. The implementation supports a single resolved phase with sub-grid bubbles governed by
@@
The Strang splitting scheme integrates the equation over one time step by integrating 3 sub-equations for $\bm{q}^{*}$, $\bm{q}^{**}$, and $\bm{q}^{***}$ as
@@
To solve for $\bm{q}^{*}$ and $\bm{q}^{***}$, where the stiff bubble source term is evaluated, we use a third-order accurate embedded Runge–Kutta scheme with adaptively controlled step sizes [72]. The step size, $h$, is updated based on the estimated error, $\bm{e}$, calculated as
@@
where $k_{1}=\partial\bm{q}^{n}/\partial t$, $k_{2}=\partial\bm{q}^{(1)}/\partial t$, and $k_{3}=\partial\bm{q}^{(2)}/\partial t$. Here, $\bm{q}^{n+1}$ represents a solution of the third-order TVD Runge–Kutta scheme, and $\hat{\bm{q}}^{n+1}$ is a 2nd-order accurate solution obtained using the same $k_{1}$, $k_{2}$, and $k_{3}$ that requires minimal additional computational cost. More details of the adaptive step size control algorithm can be found in Hairer et al. [72].
@@ 4.2.4 Low-Mach number treatment
Godunov-type Riemann solvers, including the HLLC Riemann solver, are known to lose their accuracy at low Mach numbers due to numerical dissipation in the discrete solution [73]. To address this limitation, we use two numerical treatments based on the approaches of Thornber et al. [74] and Chen et al. [75], allowing users to select their preferred scheme.
Thornber et al. [74] proposed a simple modification to the reconstructed velocities of the left and right states as
@@
On the other hand, Chen et al. [75] identified a source term, $p_{d}$, which is responsible for the wrong scaling at the low Mach number limit. To address this, they proposed an anti-dissipation pressure correction (APC) term, which is incorporated into the flux calculation. For the HLLC Riemann solver, the terms $p_{d}$ and APC are
@@ 4.2.5 HLLD Riemann Solver
For the MHD equations, the HLLD (Harten–Lax–van Leer–Discontinuities) Riemann solver [76] build on the HLL framework to resolve five of the seven characteristic waves, thereby reducing numerical dissipation and improving accuracy.
@@
where $\alpha=\mathrm{L}$ or $\mathrm{R}$. For brevity, explicit expressions for tangential velocities ($v_{\alpha}^{*}$, $w_{\alpha}^{*}$) and tangential magnetic fields ($B_{y,\alpha}^{*}$, $B_{z,\alpha}^{*}$), as well as the energy computation, are not shown here but can be found in the original paper [76].
The Alfvén wave speeds are determined by:
@@
Similarly, the corresponding expressions for $w^{**}$, $B_{y}^{**}$, and $B_{z}^{**}$ in the inner states, along with the associated energy calculation, follow a similar pattern and can be found in [76].
@@ 4.3 Software and high-performance computing
4.3.1 GPU Offloading with OpenACC on NVIDIA and AMD Hardware
Vendor portable GPU offloading on AMD and NVIDIA hardware is implemented using OpenACC [77]. OpenACC is a directive-based tool that allows the developer to generate GPU kernels by adding directive clauses around developer-identified regions of parallelism in the code. OpenACC is supported by NVIDIA’s NVHPC compiler for use with NVIDIA GPUs and HPE’s Cray Compiler Environment for use with both NVIDIA and AMD GPUs. Compilers like GNU 14 and Flang 18 support OpenACC, though their relative immaturity at time of writing makes them insufficient to support MFC.
LABEL:lst:kernel shows an example of a typical GPU kernel in MFC. The outer loops over j, k, and l iterate through $\mathcal{O}(100)$ elements each, corresponding to the grid cells in the MPI subdomain. The inner loop over i iterates through each equation in the current model. For two-phase flow problems, this loop spans $\mathcal{O}(1)$ elements. When additional features like hypo- and hyperelasticity or chemistry are added to the problem, the extent of this inner loop can be $\mathcal{O}(10)$.
OpenACC describes parallelism using terminology in which gangs, workers, and vectors refer to blocks, warps, and threads in CUDA notation. By default, when OpenACC encounters a parallel loop clause, it distributes the workload across gangs, leaving each gang to utilize a single vector. Appending the gang vector clause to the parallel loop tells the compiler to split the work across multiple gangs with a fixed vector length, which more efficiently uses available resources and increases execution speed. Appending collapse(3) to the three outer loops instructs the compiler to collapse those three loops and select the optimal gang and vector sizes based on the current problem and architecture. Our tests show that serialization of the innermost loop i increases performance, in part due to its relatively small range.
Listing 1: Directive setup for a typical MFC OpenACC kernel.
!\$acc parallel loop vector gang collapse(3) &
!\$acc default(present) private(...)
@@ 4.3.1 GPU Offloading with OpenACC on NVIDIA and AMD Hardware
! << Core kernel here, O(100) arithmetic operations >>
@@
MFC also uses NVIDIA and AMD’s vendor-provided libraries cuTensor and hipBLAS to perform hardware-optimized array reshaping for coalescing memory before the most expensive kernels. Memory coalescence before the WENO reconstruction and approximate Riemann solver kernels results in a ten-fold speedup due to the increased throughput of high-bandwidth memory. cuTENSOR [78] performs similarly to fully collapsed OpenACC loops on NVIDIA hardware, providing only marginal speedup. hipBLAS [79] performs array transposes approximately seven times faster than fully collapsed OpenACC loops. NVIDIA and CCE’s cuFFT [80] and hipFFT [81] perform fast Fourier transforms on GPU devices and FFTW [82] is used for CPU-based simulations. Code encased by OpenACC host_data use_device and end host_data clauses executed are on the GPU using these vendor-provided libraries, reducing the need for hardware-specific optimizations.
@@ 4.3.2 Performance on various CPU, GPU, and APU architectures
MFC 5.0 has been benchmarked on a broad range of CPU, GPU, and APU (superchip) devices. LABEL:t:serialperf reports the single-device performance in terms of its grind time, which is computed as time per grid point, PDE, and right-hand side evaluation. The grind times are calculated for one GPU or APU device or one CPU socket. The benchmarks were performed under a similar strategy as that used on published testbed reports [83]. These quantities are for a typical, representative compressible multi-component problem: 3D, inviscid, 5-equation model problem with two advected species (8 PDEs) and 8M grid points (158-cubed 3D uniform grid). The equation is solved via fifth-order accurate WENO finite volume reconstruction and the HLLC approximate Riemann solver. This case is in the MFC source code under examples/3D_performance_test. One can execute it via
./mfc.sh run examples/3D_performance_test/case.py -t pre_process simulation --case--optimization -n <num_cores_or_gpus> -j <num_build_threads>
This command can be executed for CPU cases. It builds an optimized version of the code for this case and then executes it. For benchmarking GPU devices, one will likely want to use -n <num_gpus> where the usual case for single GPU device benchmarking leaves <num_gpus> as unity. Similar performance is also seen for other problem configurations, such as the Euler equations (4 PDEs). All results are for the compiler that gave the best performance, including AOCC, Intel, GCC, CCE, and NVHPC.
CPU results may be performed on CPUs with more cores than reported in LABEL:t:serialperf; we report results for the best performance given the full processor die by checking the performance for different core counts on that device. CPU results are the best performance we achieved using a single socket (or die). GPU results are for a single GPU device. For single-precision (SP) GPUs, we performed computation in double-precision via conversion in compiler/software; these numbers are not for single-precision computation. AMD MI250X and MI300A devices have multiple graphics compute dies per socket; we report results for two GCDs for the AMD MI250X and the entire APU (6 XCDs) for the AMD MI300A.
Table 1: Observed grind time performance (time), reported as nanoseconds per grid point per PDE per right-hand side evaluation. Using more GPU devices or CPU sockets may not improve the grind time for the same problem, as the problem will be made smaller per marshaled device (and so grind time is expected to be approximately constant).
@@
NVIDIA GH200 APU 1 GPU 0.32 NVIDIA A10 GPU 1 GPU 4.3
NVIDIA H100 SXM5 GPU 1 GPU 0.38 AMD EPYC 7713 CPU 64 cores 5.0
NVIDIA H100 PCIe GPU 1 GPU 0.45 Intel Xeon 8480CL CPU 56 cores 5.0
@@
Intel Xeon 6740E CPU 92 cores 4.2
Figure 4 shows the roofline performance for the most expensive kernels on NVIDIA and AMD accelerators. At the time of writing, roofline analysis is unavailable on the MI300A APU, and so is not shown. NVHPC 24.5 compilers and NVIDIA Nsight Compute profilers produce roofline results on NVIDIA devices, and CCE 18 with Omniperf profile the roofline on the AMD MI250X GPU. The WENO and approximate Riemann kernels achieve 37% and 14% of peak performance on the NVIDIA A100. On the NVIDIA GH200, the WENO and approximate Riemann kernels achieve 24% and 5% of peak compute performance. The AMD MI250X achieves 22% of peak compute performance in the WENO kernel and 3% of peak performance in the approximate Riemann problem.
@@
The breakdown of time spent in the two most expensive kernels, time spent packing arrays, and time spent doing all other computations and communication is shown in fig. 5. The WENO and approximate Riemann kernels take up just over 50% of the total time on NVIDIA GPUs. On AMD GPU devices, the time spent in the WENO and approximate Riemann kernels is similar to that of NVIDIA devices. Still, these kernels account for about a third of the total wall time due to an increase in the time spent packing arrays for coalesced memory access. The profiles show a high number of cache misses on the AMD MI250X, which we expect is due to the smaller device caches than those of the shown NVIDIA device.

### v4

@@ 4.2.2 WENO-Z, TENO, and high-order non-uniform reconstruction
The WENO-Z scheme [borges2008improved] enhances classical WENO methods by improving their spectral properties while being faster than WENO-M. The (2k1)(2k-1)-th order WENO-Z weight $\omega_{r}^{(\mathrm{Z})}$ for the $r$-th stencil is computed as
@@
The Targeted Essentially Non-Oscillatory (TENO) scheme further reduces numerical dissipation by employing a stencil selection process similar to the ENO method [fu2016family]. A $K$-th order accurate TENO scheme uses K2K-2 candidate stencils of increasing width. Its equations are given by
@@
where $C_{T}$ is a smoothness threshold for stencil activation and $\tau$ is the WENO-Z parameter.
The explicit expressions for the seventh-order WENO reconstruction in our code are omitted here due to their considerable length. Instead, we present a generalized approach for reconstructing a (2k1)(2k-1)-order scheme from cell averages on non-uniform grids. For each stencil, the reconstruction polynomial is obtained using the Lagrange interpolation method [shu1997essentially]:
@@
Here, $x_{j}$ denotes cell-center location, $x_{j+1/2}$ the cell-boundary location, and $\Delta x_{j}:=x_{j+1/2}-x_{j-1/2}$ is the cell width. For each stencil $r=0,1,\dots,k-1$, the $k$-cell stencil spans the interval from $x_{-1/2}$ to $x_{k-1/2}$. Consequently, the overall (2k1)(2k-1)-cell stencil is centered at $x_{k-r-1}$, and the right flux is $p_{r}:=p(x_{k-r-1/2})$.
@@
where $\delta_{jm}$ is the Kronecker delta. The coefficients $\hat{b}_{r,j,m}$ for each cross term $\Delta f_{j}\Delta f_{m}$ and stencil $r$ are precomputed. Reformulating in terms of $\Delta f$ reduces the number of terms from $k$ to k1k-1 for $p_{r}$, and from $k(k+1)/2$ to (k1)k/2(k-1)k/2 for $\beta_{r}$, thereby reducing the computational cost.
Figure 12: Convergence results for a 1D linear advection problem with WENO-JS, WENO-Z, and TENO schemes at orders of accuracy 3, 5, and 7.
The convergence study in fig. 12 demonstrates that WENO-Z and TENO recover their intended design orders, consistent with the original WENO-JS for 3rd-, 5th-, and 7th-order reconstruction.
@@ 4.2.3 Strang splitting for stiff sub-grid dynamics
Sub-grid bubbles often exhibit rapid dynamics, necessitating a small time step for stable computation of the bubble source term. When the time scale of the background flow significantly exceeds that of the sub-grid bubbles, fully resolving the temporal dynamics becomes computationally prohibitive. To address this, we implement the Strang splitting scheme, which separates the time integration of the background flow and the sub-grid bubbles [strang1968construction]. The implementation supports a single resolved phase with sub-grid bubbles governed by
@@
The Strang splitting scheme integrates the equation over one time step by integrating three sub-equations for $\bm{q}^{*}$, $\bm{q}^{**}$, and $\bm{q}^{***}$ as
@@
To solve for $\bm{q}^{*}$ and $\bm{q}^{***}$, where the stiff bubble source term is evaluated, we use a third-order accurate embedded Runge–Kutta scheme with adaptively controlled step sizes [hairer93solving]. The step size, $h$, is updated based on the estimated error, $\bm{e}$, calculated as
@@
where $k_{1}=\partial\bm{q}^{n}/\partial t$, $k_{2}=\partial\bm{q}^{(1)}/\partial t$, and $k_{3}=\partial\bm{q}^{(2)}/\partial t$. Here, $\bm{q}^{n+1}$ represents a solution of the third-order TVD Runge–Kutta scheme, and $\hat{\bm{q}}^{n+1}$ is a 2nd-order accurate solution obtained using the same $k_{1}$, $k_{2}$, and $k_{3}$ that requires minimal additional computational cost. More details of the adaptive step size control algorithm can be found in hairer93solving.
We demonstrate the utility of the operator splitting scheme with adaptive time stepping for a 0D simulation of bubble dynamics. When a bubble is located in a high-pressure liquid, it collapses in radial size and rebounds (size increases) repeatedly, eventually damping due to viscous dissipation. Here, we test our implementation via a $\bm{u}=\bm{0}$ case and initial pressure field such that the liquid pressure is $1427$-times larger than the bubbles’ internal pressure $p_{b}$. The bubbles are assumed to be monodisperse in size and air-filled with equilibrium radius $R_{\mathrm{eq}}=$5\text{\,}\mathrm{mm}\text{/}$\$. The 0D simulations are implemented as spatially homogeneous 1D cases.
[…]
Figure 13: Evolution of bubble radius at pressure ratio $p_{\infty}/p_{b}=1427$ with and without the Strang splitting scheme and adaptive time stepping. $t_{c}$ is the total collapse time [brennen1995]. The right plot is a magnified view of the boxed region near the rebound in the left plot.
@@ 4.2.4 Low-Mach number treatment
Godunov-type Riemann solvers, including the HLLC Riemann solver, are known to lose their accuracy at low Mach numbers due to numerical dissipation in the discrete solution [guillard1999behaviour]. To address this limitation, we use two numerical treatments based on the approaches of thornber2008improved and chen2022anti, allowing users to select their preferred scheme.
thornber2008improved proposed a simple modification to the reconstructed velocities of the left and right states as
@@
On the other hand, chen2022anti identified a source term, $p_{d}$, which is responsible for the wrong scaling at the low Mach number limit. To address this, they proposed an anti-dissipation pressure correction (APC) term, which is incorporated into the flux calculation. For the HLLC Riemann solver, the terms $p_{d}$ and APC are
@@
The HLLC Riemann solver with correction schemes is validated with the 2D Gresho vortex, which is an exact solution to the incompressible Euler equations [gresho1990]. The angular velocity and pressure are given by
$u_{\phi}=u_{r}\left\{\begin{array}[]{ll}r/R,&0\leq r<R,\\ 2-r/R,&R\leq r\leq 2R,\\ 0,&2R\leq r,\end{array}\right$.
$p=p_{r}+\rho_{r}u_{r}^{2}\left\{\begin{array}[]{ll}\left(r/R\right)^{2}/2,&0\leq r<R,\\ \left(r/R\right)^{2}/2+4\left(1-\left(r/R\right)+\ln\left(r/R\right)\right),&R\leq r\leq 2R,\\ -2+4\ln 2,&2R\leq r,\end{array}\right$.
where $r=\sqrt{x^{2}+y^{2}}$ is the radial coordinate, $u_{r}=2\pi RM_{r}$ and $M_{r}=u_{r}/\sqrt{\gamma\left(p_{r}/\rho_{r}+u_{r}^{2}/2\right)}$ are reference velocity and reference Mach number, respectively. The corresponding reference time scale is $t_{r}=2\pi R/u_{r}=1/M_{r}$, which is the time for one vortex turn around. Since the velocity and pressure fields are in equilibrium, the flow is expected to be time-stationary. Figure 14 shows the local Mach number at $t/t_{r}=1$ at four reference Mach numbers: $M_{r}=10^{-1}$, $10^{-2}$, $10^{-3}$, and $10^{-4}$. The standard HLLC approximate Riemann solver shows a loss of accuracy for smaller $M_{r}$. In contrast, the HLLC solver with correction schemes preserves the flow field to visual accuracy for all tested Mach numbers.
Figure 14: Ratio of local Mach number, $M$, to reference Mach number, $M_{r}$, for the Gresho vortex at $t/t_{r}=1$ with $M_{r}=10^{-1}$, $10^{-2}$, $10^{-3}$ and $10^{-4}$: (top row) without correction; (middle row) chen2022anti; (bottom row) thornber2008improved.
@@ 4.2.5 HLLD Riemann Solver
For the MHD equations, the HLLD (Harten–Lax–van Leer–Discontinuities) Riemann solver [miyoshi2005multi] builds on the HLL framework to resolve five of the seven characteristic waves, thereby reducing numerical dissipation and improving accuracy.
@@
where $\alpha=\mathrm{L}$ or $\mathrm{R}$. For brevity, explicit expressions for tangential velocities ($v_{\alpha}^{*}$, $w_{\alpha}^{*}$) and tangential magnetic fields ($B_{y,\alpha}^{*}$, $B_{z,\alpha}^{*}$), as well as the energy computation, are not shown here but are available in miyoshi2005multi. The Alfvén wave speeds are determined by:
@@
The corresponding expressions for $w^{**}$, $B_{y}^{**}$, and $B_{z}^{**}$ in the inner states, along with the associated energy calculation, follow a similar pattern and can be found in miyoshi2005multi.
@@
To validate the HLLD implementation, we use the standard one-dimensional MHD Brio–Wu shock-tube problem from miyoshi2005multi. The density profile in Figure 10 (a) closely matches the reference, as do the other variables (not shown).
@@ 4.3 Software and high-performance computing
4.3.1 APU and GPU offloading with OpenACC and OpenMP
Table 1: Compiler support for GPU offload to Nvidia hardware, AMD hardware, and for unified shared memory mode on AMD hardware. NVHPC SDK 25.9, CCE 19.0.0, and AMD Flang Preview 7.0.5 are used to populate this table. Support for both the AMD and CCE compilers is provided to enable AMD GPU offloading when CCE is not available.
OpenMP OpenACC
Compiler NVIDIA GPU AMD A/GPU AMD USM NVIDIA GPU AMD A/GPU AMD USM
AMD
CCE
NVHPC
Vendor portable GPU offloading on AMD and NVIDIA hardware is implemented using OpenACC [Wienke2012] and OpenMP [OpenMP]. OpenACC and OpenMP are directive-based tools that enable developers to generate GPU kernels by adding directive clauses around regions of parallelism identified in the code. Table 1 summarizes the current state of compiler support for OpenMP and OpenACC on NVIDIA and AMD GPUs, and for unified shared memory mode on AMD hardware. OpenACC is the preferred offloading tool in MFC due to its superior performance over OpenMP on NVIDIA and AMD hardware [wilfongGB25]. OpenMP is used as a fallback when unified shared memory (USM) is required on AMD hardware. Support for AMD compilers is provided for using AMD GPUs when CCE is not available. Compilers like GNU 14 and Flang 18 support OpenACC, though their relative immaturity at the time of writing makes them insufficient to support MFC.
Support for OpenACC and OpenMP is implemented through a single source code version using the Fortran preprocessor Fypp [fypp] (additional uses of Fypp in MFC are described in section 4.3.6). LABEL:lst:kernel shows an example of a typical GPU kernel in MFC. Parallel regions are wrapped with the GPU_PARALLEL_LOOP and END_GPU_PARALLEL_LOOP macros, which expand to OpenACC or OpenMP directives depending on the selected offloading tool. The outer loops over j, k, and l iterate through $\mathcal{O}(100)$ elements each, corresponding to the grid cells in the MPI subdomain. The inner loop over i iterates through each equation in the current model. For two-phase flow problems, this loop spans $\mathcal{O}(1)$ elements. When additional features like hypo- and hyperelasticity or chemistry are added to the problem, the extent of this inner loop can be $\mathcal{O}(10)$.
Listing 1: Directive setup for a typical MFC OpenACC kernel. The Fypp macros GPU_PARALLEL_LOOP and END_GPU_PARALLEL_LOOP expand to OpenACC or OpenMP directives depending on the selected offloading tool.
\$:GPU_PARALLEL_LOOP(collapse=3, private=’[...]’)
@@ 4.3.1 APU and GPU offloading with OpenACC and OpenMP
\$:GPU_LOOP(parallelism=’[seq]’)
do i = 1, num_PDEs
! Core kernel here
! O(100) arithmetic operations
end do
end do
end do
end do
\$:END_GPU_PARALLEL_LOOP
LABEL:lst:kernelACC shows the OpenACC directives generated by the Fypp macros in LABEL:lst:kernel. OpenACC describes parallelism using terminology in which gangs, workers, and vectors refer to blocks, warps, and threads in CUDA notation. By default, when OpenACC encounters a parallel loop clause, it distributes the workload across gangs, leaving each gang to utilize a single vector. Appending the gang vector clause to the parallel loop tells the compiler to split the work across multiple gangs with a fixed vector length, which more efficiently uses available resources and increases execution speed. Appending collapse(3) to the three outer loops instructs the compiler to collapse those three loops and select the optimal gang and vector sizes based on the current problem and architecture. Our tests show that serialization of the innermost loop i increases performance, in part due to its relatively small range.
LABEL:lst:kernelMP shows the OpenMP directives generated for the same kernel by the Fypp macros in LABEL:lst:kernel. OpenMP offloads parallel regions to the GPU when it encounters an omp target directive. Adding teams after omp target instructs the compiler to launch a league of teams, which typically map to CUDA thread blocks on NVIDIA GPUs. The combined teams distribute parallel do construct then further decomposes the iteration space: distribute partitions iterations across teams (i.e., across blocks), and parallel do partitions iterations across threads within each team. The defaultmap clauses specify how scalars, aggregates, allocatables, and pointers should be implicitly mapped to the device, since OpenMP requires more explicit data-mapping semantics than OpenACC. Adding collapse(3) collapses the three nested loops into a single iteration space, improving load balance and work distribution across teams and threads, just as in OpenACC.
Listing 2a: OpenACC directives generated by Fypp macros.
!\$acc parallel loop vector gang collapse(3) &
!\$acc default(present) private(...)
do l = 0, p ! Z-direction
do k = 0, n ! Y-direction
do j = 0, m ! X-direction
@@
! Core kernel
! O(100) - O(1000) operations
@@
Listing 2b: OpenMP directives generated by Fypp macros.
!\$omp target teams distribute parallel do &
!\$omp simd defaultmap(firstprivate:scalar) &
!\$omp defaultmap(tofrom:aggregate) &
!\$omp defaultmap(present:allocatable) &
!\$omp defaultmap(present:pointer) &
!\$omp collapse(3) private(...)
do l = 0, p ! Z-direction
do k = 0, n ! Y-direction
do j = 0, m ! X-direction
!\$omp loop bind(thread)
do i = 1, num_PDEs
! Core kernel
! O(100) - O(1000) operations
end do
end do
end do
end do
!\$omp end target teams distribute parallel do
MFC also uses NVIDIA and AMD’s vendor-provided libraries cuTensor and hipBLAS to perform hardware-optimized array reshaping for coalescing memory before the most expensive kernels. Memory coalescence before the WENO reconstruction and approximate Riemann solver kernels results in a ten-fold speedup due to the increased throughput of high-bandwidth memory (HBM). cuTENSOR [cuTENSOR] performs similarly to fully collapsed OpenACC loops on NVIDIA hardware, providing only marginal speedup. hipBLAS [hipBLAS] performs array transposes approximately seven times faster than fully collapsed OpenACC loops. NVIDIA and CCE’s cuFFT [cuFFT] and hipFFT [hipFFT] perform fast Fourier transforms on GPU devices, and FFTW [FFTW] is used for CPU-based simulations. Code enclosed by OpenACC host_data use_device and end host_data or OpenMP target data and end target data clauses are executed on the GPU using these vendor-provided libraries, reducing the need for hardware-specific optimizations.
@@ 4.3.2 Performance on various CPU, GPU, and APU architectures
MFC 5.0 has been benchmarked on a broad range of CPUs, GPUs, and APUs (also known as superchips). LABEL:t:serialperf reports the single-device performance in terms of its grind time, which is computed as time per grid point, PDE, and right-hand side evaluation. The grind times are calculated for one GPU, APU, or CPU socket. The benchmarks were performed using a similar strategy to that employed in published testbed reports [elwasif]. These quantities are for a typical, representative compressible multi-component problem: 3D, inviscid, 5-equation model problem with two advected species (8 PDEs) and 8M grid points (158-cubed 3D uniform grid). The equation is solved via fifth-order accurate WENO finite volume reconstruction and the HLLC approximate Riemann solver. This case is in the MFC source code under examples/3D_performance_test. One can execute it via
./mfc.sh run examples/3D_performance_test/case.py -t pre_process simulation --case-optimization -n <num_cores_or_gpus> -j <num_build_threads>
This command can be executed for CPU cases. It builds an optimized version of the code for this case and then executes it. For benchmarking GPU devices, one will likely want to use -n <num_gpus> where the usual case for single GPU device benchmarking leaves <num_gpus> as unity. Similar performance is also observed for other problem configurations, such as the Euler equations (4 PDEs). All results are for the compiler that gave the best performance, including AOCC, Intel, GCC, CCE, and NVHPC.
CPU results may be performed on CPUs with more cores than reported in LABEL:t:serialperf; we report results for the best performance, given the full processor die, by checking performance for different core counts on that device. CPU results are the best performance we achieved using a single socket (or die). GPU results are for a single GPU device. For single-precision (SP) GPUs, we performed computations in double-precision via compiler/software conversion; these numbers are not representative of single-precision computations. AMD MI250X and MI300A devices have multiple graphics compute dies per socket; we report results for two GCDs for the AMD MI250X and the entire APU (6 XCDs) for the AMD MI300A.
Table 2: *
@@
NVIDIA GH200 APU 1 GPU 0.32 Intel Xeon 6740E CPU 92 cores 4.2
NVIDIA H100 SXM5 GPU 1 GPU 0.38 NVIDIA A10 GPU 1 GPU 4.3
NVIDIA H100 PCIe GPU 1 GPU 0.45 AMD EPYC 7713 CPU 64 cores 5.0
NVIDIA B200 GPU 1 GPU 0.46 Intel Xeon 8480CL CPU 56 cores 5.0
@@
Cross-module inlining not yet supported.
Figure 15 shows the roofline performance for the most expensive kernels on NVIDIA and AMD accelerators. At the time of writing, roofline analysis is not available on the MI300A APU and therefore is not shown. NVHPC 24.5 compilers and NVIDIA Nsight Compute profilers produce roofline results on NVIDIA devices, and CCE 18 with Omniperf profiles of the roofline on the AMD MI250X GPU. The WENO and approximate Riemann kernels achieve 37% and 14% of peak performance on the NVIDIA A100. On the NVIDIA GH200, the WENO and approximate Riemann kernels achieve 24% and 5% of peak compute performance. The AMD MI250X achieves 22% of peak compute performance in the WENO kernel and 3% of peak performance in the approximate Riemann problem.
@@
The breakdown of time spent in the two most expensive kernels, time spent packing arrays, and time spent doing all other computations and communication is shown in fig. 16. The WENO and approximate Riemann kernels account for about 50% of the total wall time on NVIDIA GPUs. On AMD GPU devices, the time spent in the WENO and approximate Riemann kernels is similar to that of NVIDIA devices. Still, these kernels account for about a third of the total wall time due to an increase in the time spent packing arrays for coalesced memory access. The profiles show a high number of cache misses on the AMD MI250X, which we expect is due to the smaller device caches compared to those of the shown NVIDIA device.